OpenStudy (anonymous):

Which of the following is a factor of x2+ 2x − 48? x − 8 x + 12 x − 6 None of the above

Nnesha (nnesha):

what two number you multiply to get -48 or when you add or subtract them you should get middle term(2) easy to factor it AC method or headphone method find two number when you multiply them you should get product of AC and when you add or subtract the;m you should get middle term b

OpenStudy (anonymous):

ok im confused i know its either x-8 or x-6

Nnesha (nnesha):

-6 times -8 = positive 48 (we need negative 48) -6-8 = -14 so one of them suppose to be positive you tell which one should be negative ?

OpenStudy (anonymous):

x-8?

Nnesha (nnesha):

oky check it ot -8 times 6 = -48 correct but -8+6 = positive 2 ?

OpenStudy (anonymous):

oh crap okay so it's x-6

Nnesha (nnesha):

yes right -6+8 = +2